# A Study of the Framework and Real-World Applications of Language Embedding for 3D Scene Understanding

### v1

@@ 7.1. Human Reconstruction and Avatar Generation
Modeling humans in 3D is a vital task across multiple industries, including entertainment, gaming, fashion, and e-commerce. Whether for digital avatars in virtual worlds, realistic virtual try-on systems, or expressive animated characters, accurately capturing the complexity of human appearance, motion, and expression remains a significant challenge. Gaussian splatting has emerged as a powerful technique for 3D reconstruction, offering a fast and efficient way to generate highly detailed human models from real-world data. However, beyond static reconstructions, there is increasing interest in integrating language to enhance customization, interactivity, and ease of creation. This section explores how 3DGS is applied to human modeling, how diffusion models contribute by generating multi-view images for improved reconstructions, and how language-based interfaces enable more accessible and flexible 3D generation.
Human Priors and Parametric Models: One key approach to human reconstruction relies on parametric models like SMPL (Loper et al., 2023; Pavlakos et al., 2019) and imGHUM (Alldieck et al., 2021), which provide a canonical human body shape that can be easily adjusted for variations in pose, gender, and body type. These models have been widely adopted due to their flexibility in representing diverse human forms, enabling detailed manipulations such as animation and shape adjustments. While such models provide a solid foundation for representing human structure, they still struggle to capture fine-grained surface details, such as skin textures or clothing dynamics, leaving room for improvement through the combination of more advanced rendering methods such as NeRF and 3DGS (Yang et al., 2024).
Gaussian splatting, as a method of 3D reconstruction, has gained attention for its ability to represent complex organic shapes, like the human body, with high fidelity and computational efficiency. Unlike traditional mesh-based models, which rely on polygons and vertices, 3DGS allows for smoother surfaces and transitions, making it particularly well-suited for capturing the fluidity of human movement and skin details, such as wrinkles or folds in clothing, which are often lost in traditional approaches. Furthermore, by relying on point cloud representations, Gaussian splatting enables more natural and flexible reconstructions, making it ideal for creating lifelike avatars in dynamic environments.
Recent works have sought to combine the strengths of SMPL-based priors with 3DGS to overcome the limitations of each individual technique (Liu et al., 2024c; Moon et al., 2024; Gong et al., 2024; Kocabas et al., 2024). By using SMPL as the parametric structure for the human body and applying 3DGS for surface details, this hybrid approach enables more realistic reconstructions. 3DGS not only refines the appearance of the body but also enhances the fluidity of animation, making it easier to generate dynamic poses or modify expressions in real time. This combination is particularly useful for applications such as virtual avatars or digital humans, where precise control over body structure and appearance is essential.
@@ 7.2. Immersive Technology
The initial taxonomy of immersive technologies was presented by Milgram et al. in (Milgram et al., 1995), conceptualizing different levels of immersion through the reality-virtuality continuum. This framework attempted to define how then emerging technologies bridge the gap between the real and virtual worlds. In this paper, Extended Reality (XR) is used as an encompassing term for Virtual Reality (VR), Augmented Reality (AR), and Mixed Reality (MR). The varying degrees of immersion enabled by XR technologies provide a spectrum of human-technology interaction interfaces, tailored to both digital and physical environments. Today, XR is widely used in entertainment, education, robotics, cyber-physical systems, healthcare, and other domains (Adriana Cárdenas-Robledo et al., 2022).
Creating interactive 3D environments for XR applications is often time-consuming and resource-intensive. While traditional methods require extensive data capture and processing, recent advances like 3DGS offer an effective, real-time solution for high-fidelity scene reconstruction, making it well-suited for immersive VR and AR experiences where responsiveness and visual quality are essential.
3DGS in Extended Reality: 3DGS offers real-time, high-fidelity scene representation, making it a strong candidate for XR applications. Its compact, explicit representation supports photorealistic rendering, enhancing immersion and reducing cybersickness. However, challenges remain around physical interaction, real-time performance, and integration with XR development tools.
@@
Despite these advances, core limitations persist. 3DGS lacks an underlying mesh structure, hindering collision detection and physics-based interaction essential for most XR SDKs. Mesh extraction methods (Guédon and Lepetit, 2024; Waczynska et al., 2024; Wolf et al., 2024) often produce dense, noisy outputs. Similarly, material properties and lighting behaviors require manual tuning or ML models, which have not yet been adapted for XR contexts. For example, Gao et al. (2024); Moenne-Loccoz et al. (2024) explore relighting and reflections in 3DGS, but lack XR integration. A promising direction is the hybrid mesh-splat representation proposed by Dongye et al. (2025), which dynamically selects mesh or splat representations based on real-time user intent.
@@ 8. Challenges and Future Research Directions
Despite recent progress in language-embedded 3D Gaussian Splatting, several fundamental challenges remain. These span architectural, computational, and dataset-related limitations, as well as unresolved issues in semantic fidelity, spatial reasoning, and generalization. In this section, we highlight key open research problems and outline promising directions to address them, with a focus on improving scalability, robustness, and semantic alignment across real-world scenarios.
8.1. Photorealism vs. Semantic Fidelity in 3D Gaussian Splatting

### v2

@@ 7.1. Human Reconstruction and Avatar Generation
Modeling humans in 3D is a vital task across multiple industries, including entertainment, gaming, fashion, and e-commerce. Whether for digital avatars in virtual worlds, realistic virtual try-on systems, or expressive animated characters, accurately capturing the complexity of human appearance, motion, and expression remains a significant challenge. While mesh-based and parametric models (such as SMPL (Loper et al., 2023; Pavlakos et al., 2019) and imGHUM (Alldieck et al., 2021)) provide a structured representation of body shape and pose, they often fail to capture fine-grained details such as facial expressions, skin texture, or clothing deformation (Yang et al., 2024). Recent works address these limitations by combining 3DGS with parametric priors and language-based editing, enabling high-quality, customizable human avatars (Liu et al., 2024c; Moon et al., 2024; Gong et al., 2024; Kocabas et al., 2024).
Several methods incorporate diffusion-based optimization guided by text prompts to generate 3D avatars with realistic appearance and geometry (Liu et al., 2024c; Gong et al., 2024; Yuan et al., 2024). Others focus on controllable head avatars, stylization, or emotion-driven animation using vision-language models such as CLIP or BLIP-2 (Zhou et al., 2024b; Liu et al., 2024a; Jiang et al., 2024a). For applications like virtual try-on, hybrid pipelines leverage latent diffusion, ControlNet, or LoRA-based personalization to support garment-specific edits while maintaining multi-view consistency (Chen et al., 2024b; Cao et al., 2024).
Together, these systems demonstrate the value of combining 3DGS with foundation models to enable expressive, dynamic, and user-controllable human reconstruction.
@@ 7.2. Immersive Technology
Immersive technologies such as Virtual Reality (VR), Augmented Reality (AR), and Mixed Reality (MR), collectively referred to as Extended Reality (XR), enable rich, real-time interaction with 3D environments (Adriana Cárdenas-Robledo et al., 2022). When combined with language-embedded 3DGS, these systems support high-fidelity reconstruction, semantic understanding, and intuitive user interaction across a range of XR applications.
3DGS in Extended Reality: 3DGS offers real-time, high-fidelity scene representation, making it a strong candidate for XR applications. Its compact and explicit representation supports photorealistic rendering, contributing to immersive virtual environments. However, challenges remain around physical interaction, real-time performance, and integration with XR development tools.
@@
While integration with physics engines and XR SDKs remains an open challenge, ongoing work, including hybrid mesh-splat models (Dongye et al., 2025), suggests promising pathways toward more interactive and physically grounded XR experiences.
@@ 8. Challenges and Future Research Directions
Despite recent progress in language-embedded 3DGS, several fundamental challenges remain. These span architectural, computational, and dataset-related limitations, as well as unresolved issues in semantic fidelity, spatial reasoning, and generalization. In this section, we highlight key open research problems and outline promising directions to address them, with a focus on improving scalability, robustness, and semantic alignment across real-world scenarios.
8.1. Photorealism vs. Semantic Fidelity in 3DGS